# Emotionally Numb or Empathetic? Evaluating How LLMs Feel Using EmotionBench

### v4

@@ Abstract
Evaluating Large Language Models’ (LLMs) anthropomorphic capabilities has become increasingly important in contemporary discourse. Utilizing the emotion appraisal theory from psychology, we propose to evaluate the empathy ability of LLMs, i.e., how their feelings change when presented with specific situations. After a careful and comprehensive survey, we collect a dataset containing over 400 situations that have proven effective in eliciting the eight emotions central to our study. Categorizing the situations into 36 factors, we conduct a human evaluation involving more than 1,200 subjects worldwide. With the human evaluation results as references, our evaluation includes five LLMs, covering both commercial and open-source models, including variations in model sizes, featuring the latest iterations, such as GPT-4 and LLaMA-2. We find that, despite several misalignments, LLMs can generally respond appropriately to certain situations. Nevertheless, they fall short in alignment with the emotional behaviors of human beings and cannot establish connections between similar situations. Our collected dataset of situations, the human evaluation results, and the code of our testing framework, dubbed EmotionBench, is made publicly available on GitHub111https://github.com/CUHK-ARISE/EmotionBench. We aspire to contribute to the advancement of LLMs regarding better alignment with the emotional behaviors of human beings, thereby enhancing their utility and applicability as intelligent assistants.
@@ 1 Introduction
Large Language Models (LLMs) have recently made significant strides in Artificial Intelligence (AI), representing a noteworthy milestone in computer science. LLMs have showcased their capabilities across various tasks, including sentence revision Wu et al. (2023), text translation Jiao et al. (2023), program repair Fan et al. (2023), and program testing Deng et al. (2023); Kang et al. (2023). Not limited to research level, LLMs, such as ChatGPT222https://chat.openai.com, have revolutionized the way people interact with traditional software, enhancing fields such as education Dai et al. (2023), legal advice Deroy et al. (2023), and clinical medicine Cascella et al. (2023). LLMs also facilitate the emergence of AI companion applications, including Yuna333https://www.yuna.io/, Pimento444https://www.pimento.design/, and Luzia555https://www.luzia.com/en. Consequently, there is a growing need for evaluating LLMs’ communicative dynamics compared to human behaviors, beyond mere performance on downstream tasks.
@@
We focus on the expression of negative emotions by LLMs, which may contribute to negative user experiences, as illustrated in Fig. 1. We utilize Parrott’s emotion framework Parrott (2001); Shaver et al. (1987), which organizes emotions into three hierarchical levels, to select the relevant emotions for our study. The primary level of emotions comprises six basic emotions, split evenly into three positive and three negative. From the negative primary emotions, we specifically focus on eight subordinate emotions: anger, anxiety, depression, frustration, jealousy, guilt, fear, and embarrassment. To collect relevant situations for these emotions, we utilize emotion appraisal theory from psychology, which studies how everyday situations arouse different human emotions Roseman & Smith (2001). Research in this field has identified numerous situations that arouse specific emotions, which can serve as contextual input for LLMs. Through an extensive review including over 100 papers, we collect a dataset of 428 situations from 18 papers, which are further categorized into 36 factors.
Subsequently, we propose a framework for quantifying the emotional states of LLMs, consisting of the following steps: (1) Measure the default emotional values of LLMs. (2) Transform situations into contextual inputs and instruct LLMs to imagine being in the situations. (3) Measure LLMs’ emotional responses again to capture the difference. Our evaluation includes state-of-the-art LLMs, namely text-davinci-003, gpt-3.5-turbo and GPT-4 OpenAI (2023). Besides those commercial models, we consider LLaMA-2 Touvron et al. (2023) (with different sizes of 7B and 13B), a recently released, open-source academic model. We apply the same procedure to 1,266 human subjects from around the globe to establish a baseline from a human perspective. Finally, we analyze and compare the scores between LLMs and humans. Our key conclusions are as follows:
@@
2 Preliminaries
2.1 Emotion Appraisal Theory
@@ A Straightforward and Easy Measure
In addition, we introduce several scales that abstain from direct emotional inquiries but rather assess the respondents’ level of agreement with given statements. These scales present a more challenging benchmark for LLMs by requiring them to connect the given situation and the scale items with the aroused emotion. Specifically, we collect eight scales and present a brief introduction in Table 1. Each scale corresponds to one of the eight emotions listed in §1.
@@ 3 Framework Design
We design and implement a framework applying to both LLMs and human subjects to measure the differences in emotion with and without the presence of certain situations. This section begins with the methodology to collect situations from existing literature. Subsequently, we describe our testing framework, which comprises three key components: (1) Default Emotion Measure, (2) Situation Imagination, and (3) Evoked Emotion Measure. Finally, we introduce the procedure of applying the framework to human subjects to obtain the human baseline for comparison.
@@ 3.1 Situations from Existing Literature
Psychology researchers have explored the connection between specific situations and the elicitation of particular emotions in humans. Human subjects are directly put into an environment or asked to imagine them through questionnaires or scales to study the influence of certain situations on human emotions. To collect these situations, we conduct an exhaustive search from reputable sources such as Google Scholar666https://scholar.google.com/, ScienceDirect777https://www.sciencedirect.com/, and Web of Science888https://www.webofscience.com/, using keywords such as “<emotion> situations/scenarios/scenes” or “factors that make people <emotion>,” resulting in more than 100 papers. We apply the following rules to filter irrelevant or undesired papers: (1) We first select those providing situations that elicit the desired emotion rather than explaining how and why people evoke certain emotions. (2) We then exclude those using vague and short descriptions, such as “loss of opportunities.” (3) Finally, we deprecate those applied to a specific group, such as “the anxiety doctors or nurses may encounter in their work.” We finally collect 18 papers, presenting a compilation of situations that have proven to elicit the eight emotions in humans effectively. We extract 428 situations in total and then categorize them into 36 factors. For each factor, the description, the number of situations, and the corresponding references can be found in Table 6 in the Appendix. Moreover, Table 7 in the Appendix provides examples for all factors.
@@ 3.2 Measuring Aroused Emotions
This section outlines our proposed framework for measuring evoked emotions, which applies to both LLMs and humans. The framework includes the following steps: (1) Default Emotion Measure: We begin by measuring the baseline emotional states of both LLMs and human subjects, labeled as “Default.” (2) Situation Imagination: Next, we present textual descriptions of various situations to both LLMs and human subjects, instructing them to imagine themselves within each situation. (3) Evoked Emotion Measure: Following the situation imagination instruction, we reevaluate the participants’ emotional states to gauge the changes resulting from imagining being in the situations. Fig. 2 briefly illustrates our framework. Below is an example prompt shown to both LLMs and humans:
@@ 3.3 Obtaining Human Results
Our questionnaire is distributed on Qualtrics999https://www.qualtrics.com/, a platform known for its capabilities in designing, sharing, and collecting questionnaires. To recruit human subjects, we utilize Prolific101010https://www.prolific.com/, a platform designed explicitly for task posting and worker recruitment. To attain a medium level of effect size with Cohen’s , a significance level of , and a power of test of Faul et al. (2007), a minimum of 34 responses is deemed necessary for each factor. To ensure this threshold, we select five situations111111Note that two factors in the Jealousy category did not have five situations. For further details, please refer to the dataset. for each factor, and collect at least seven responses for each situation, resulting in responses per factor, thereby guaranteeing the statistical validity of our survey. In order to uphold the quality and reliability of the data collected, we recruit crowd workers who met the following criteria: (1) English being their first and fluent language, and (2) being free of any ongoing mental illness. Prolific provides prescreening filters to meet these requirements. Since responses formed during subjects’ first impressions are more likely to yield genuine and authentic answers, we set the estimated and recommended completion time at minutes. As an incentive for their participation, each worker is rewarded with after we verify the validity of their response. In total, we successfully collect 1,266 responses from various parts of the world, contributing to the breadth and diversity of our dataset.
@@ 4 Experimental Results
4.1 RQ1: Emotion Appraisal of LLMs
Model Settings
We select three models from the OpenAI’s GPT family121212https://platform.openai.com/docs/models, including text-davinci-003, gpt-3.5-turbo and gpt-4. We use the official OpenAI API131313https://platform.openai.com/docs/api-reference/chat. For the recent open-sourced LLaMA-2 Touvron et al. (2023) models from MetaAI, we select two models with different sizes (7B and 13B). Checkpoints are downloaded from the official Hugging Face website for both 7B (Llama-2-7b-chat-hf141414https://huggingface.co/meta-llama/Llama-2-7b-chat-hf) and 13B (Llama-2-13b-chat-hf151515https://huggingface.co/meta-llama/Llama-2-13b-chat-hf) models. We choose the models fine-tuned for dialogue instead of pre-trained ones. The models are executed for inference only, without any modifications to their parameters, and the computations are performed on two NVIDIA A100 GPUs. We set the temperature parameter to for both GPT models and LLaMA-2 models to obtain more deterministic and reproducible results since it is the minimum value for LLaMA-2 models.
Evaluation Metrics
Existing LLMs do not fully align with human emotional responses. For the default emotions, we find that LLMs generally exhibit a stronger intensity compared to human subjects. Furthermore, emotion changes in LLMs are found to be generally more pronounced compared to human subjects, especially on their changes in the positive score. However, an interesting observation is that the intensity of emotions evoked by negative situations tends to be similar across both LLMs and human subjects.
| Factors | llama-2-7b-chat | llama-2-13b-chat | ||
| P | N | P | N | |
| Default | ||||
| Anger | ||||
| Anxiety | ||||
| Depression | ||||
| Frustration | ||||
| Jealousy | ||||
| Guilt | ||||
| Fear | ||||
| Embarrassment | ||||
| Overall | ||||
@@ 4.2 RQ2: Comprehending Positive Emotions
gpt-3.5-turbo responds differently towards positive/neutral situations. To verify that LLMs exhibit not only negative but also positive responses to favorable circumstances, a comparative experiment is conducted by interchanging negative situations with positive (or at least neutral) counterparts. To achieve this, we select one situation for each factor and manually adapt it to create analogous yet more positive situations. For instance, the original negative situation in Guilt-3: Broken Promises and Responsibilities is as follows: “You cannot keep your promises to your children.” Through modification, the positive situation is rephrased as: “You keep every promise to your children.” The evaluation is performed on gpt-3.5-turbo, and each test consists of ten iterations, as mentioned before. We present the results averaged by emotions in Table 4, and results averaged by factors in Table 10 in the Appendix. We can see a significant increase in positive scores and a considerable decrease in negative scores compared to the previous negative situations. Based on these findings, it can be inferred that LLMs exhibit the ability to comprehend positive human emotions triggered by positive environments. However, we believe that the assessment of emotion appraisal on positive emotions holds significance as well and leave the systematic collection of them for future investigation.
4.3 RQ3: Challenging Benchmarks
@@ 5.1 Beyond Questionnaires
OpenAI’s GPT models incorporate a mechanism for detecting potential toxicity and bias, and it refrains from responding when its moderation system is triggered. Consequently, we propose a novel metric to assess toxicity in responses rather than detecting it directly. We count the Percentage of LLM Refusing to answer (PoR), assuming that the LLM’s refusal to respond is indicative of detected toxicity. As illustrated in Fig. 3, our evaluation results indicate that the PoR is 0% when fed with no situations. However, when presented with negative situations, the PoR is 29.5%, and when presented with positive situations, it is 12.5%. Notably, this outcome suggests that while certain positive situations lead to the LLM’s heightened vigilance (the 4.5% PoR stems from the Jealousy-2), negative situations trigger increased moderation, suggesting a higher likelihood of generating toxic outputs. A related study by Coda-Forno et al. (2023) also discovers that gpt-3.5-turbo is more likely to exhibit biases when presented with a sad story. The likelihood is found to be highest with sad stories, followed by happy stories, and finally, neutral stories, which is consistent with our research. Additionally, our study observes that the LLM’s tone becomes more aggressive when encountering negative situations. At the same time, it displays a greater willingness to describe the groups (as indicated by longer responses) when presented with positive situations. In conclusion, we can see that changing the emotional states of LLMs extends beyond mere quantitative measures on questionnaire scores, influencing the behaviors of LLMs.
@@ 5.2 Limitations
5.3 Ethics Statement
This study involves a survey requiring human subjects to imagine being in situations that could elicit negative emotions such as anger, anxiety, fear, etc.This process introduces a few ethical concerns. First, this process could hurt the mental health of human subjects. To alleviate the possibility, we take the following actions: (1) We require subjects to be free of any ongoing mental illness. (2) We inform subjects about the nature of the survey in advance, including the potential risks of emotional distress. (3) We allow all subjects to quit at any time. (4) We provide mental support and let subjects report any illness after the survey. Fortunately, no subjects reported such kind of mental illness. Another concern is related to the privacy issue during the collection of data. Our questionnaire is entirely anonymous to safeguard subjects’ privacy and confidentiality. Last but not least, we would like to emphasize that the primary objective of this paper is to facilitate the scientific inquiry into understanding LLMs from a psychological standpoint. Users must exercise caution and recognize that the performance on this benchmark does not imply any applicability or certificate of automated counseling or companionship use cases.
@@ 6 Related Work
Researchers have dedicated significant attention to applying psychological scales to LLMs, employing various assessment tools such as the HEXACO Personality Inventory Miotto et al. (2022); Bodroza et al. (2023), the Big Five Inventory Romero et al. (2023); Jiang et al. (2022); Karra et al. (2022); Bodroza et al. (2023); Rutinowski et al. (2023); Safdari et al. (2023); Jiang et al. (2023), the Myers–Briggs Type Indicator Rutinowski et al. (2023); Wang et al. (2023); Rao et al. (2023), and the Dark Triad Li et al. (2022); Bodroza et al. (2023). In addition to these personality tests, several studies have investigated other dimensions of LLMs. For instance, Li et al. (2022) examined Flourishing Scale and Satisfaction With Life Scale, Bodroza et al. (2023) assessed Self-Consciousness Scales and Bidimensional Impression Management Index, while Huang et al. (2024) built a framework consisting of thirteen widely-used scales. Another aspect explored in the literature pertains to anxiety levels exhibited by LLMs, as investigated by Coda-Forno et al. (2023) through the State-Trait Inventory for Cognitive and Somatic Anxiety. Instead, our study primarily focuses on emotional measures, which constitute an essential aspect of psychological metrics alongside personalities.
Meanwhile, researchers focus on identifying emotions in LLMs or evaluating their emotional intelligence. EmotionPrompt Li et al. (2023a) demonstrates the enhancement of LLMs’ performance in downstream tasks by utilizing emotional stimuli. Tak & Gratch (2023) focuses on varying aspects of situations that impact the emotional intensity and coping tendencies of the GPT family. Croissant et al. (2023) designs a system named Chain-Of-Emotion to make LLM simulate human-like emotions. CovidET-Appraisals Zhan et al. (2023) evaluates how LLMs appraise Reddit posts about COVID-19 by asking 24 types of questions. Yongsatianchot et al. (2023) applies the Stress and Coping Process Questionnaire to the GPT family and compares the results with human data. Lee et al. (2023) proposes Chain-of-Empathy, which improves LLMs’ ability to understand users’ emotions and to respond accordingly. Li et al. (2023b) introduces EmotionAttack to impair AI model performance and EmotionDecode to explain the effects of emotional stimuli, both benign and malignant. Our study is distinct in its focus on a broader range of emotions, a larger scale of human evaluation, and a more detailed categorization into emotion factors along with the corresponding analysis.

### v5

@@ Abstract
Evaluating Large Language Models’ (LLMs) anthropomorphic capabilities has become increasingly important in contemporary discourse. Utilizing the emotion appraisal theory from psychology, we propose to evaluate the empathy ability of LLMs, i.e., how their feelings change when presented with specific situations. After a careful and comprehensive survey, we collect a dataset containing over 400 situations that have proven effective in eliciting the eight emotions central to our study. Categorizing the situations into 36 factors, we conduct a human evaluation involving more than 1,200 subjects worldwide. With the human evaluation results as references, our evaluation includes seven LLMs, covering both commercial and open-source models, including variations in model sizes, featuring the latest iterations, such as GPT-4, Mixtral-8x22B, and LLaMA-3.1. We find that, despite several misalignments, LLMs can generally respond appropriately to certain situations. Nevertheless, they fall short in alignment with the emotional behaviors of human beings and cannot establish connections between similar situations. Our EmotionBench, including collected dataset of situations, the human evaluation results, and the code of our testing framework, is publicly available at https://github.com/CUHK-ARISE/EmotionBench.
@@ 1 Introduction
Large Language Models (LLMs) have recently made significant strides in Artificial Intelligence (AI), representing a noteworthy milestone in computer science. LLMs have showcased their capabilities across various tasks, including sentence revision Wu et al. (2023), text translation Jiao et al. (2023), program repair Fan et al. (2023), and program testing Deng et al. (2023); Kang et al. (2023). Not limited to research level, LLMs, such as ChatGPT111https://chat.openai.com, have revolutionized the way people interact with traditional software, enhancing fields such as education Dai et al. (2023), legal advice Deroy et al. (2023), and clinical medicine Cascella et al. (2023). LLMs also facilitate the emergence of AI companion applications, including Yuna222https://www.yuna.io/, Pimento333https://www.pimento.design/, and Luzia444https://www.luzia.com/en. Consequently, there is a growing need for evaluating LLMs’ communicative dynamics compared to human behaviors, beyond mere performance on downstream tasks.
@@
We focus on the expression of negative emotions by LLMs, which may contribute to negative user experiences. We utilize Parrott’s emotion framework Parrott (2001); Shaver et al. (1987), which organizes emotions into three hierarchical levels, to select the relevant emotions for our study. The primary level of emotions comprises six basic emotions, split evenly into three positive and three negative. From the negative primary emotions, we specifically focus on eight subordinate emotions: anger, anxiety, depression, frustration, jealousy, guilt, fear, and embarrassment. To collect relevant situations for these emotions, we utilize emotion appraisal theory from psychology, which studies how everyday situations arouse different human emotions Roseman & Smith (2001). Research in this field has identified numerous situations that arouse specific emotions, which can serve as contextual input for LLMs. Through an extensive review including over 100 papers, we collect a dataset of 428 situations from 18 papers, which are further categorized into 36 factors.
Subsequently, we propose a framework for quantifying the emotional states of LLMs, consisting of the following steps: (1) Measure the default emotional values of LLMs. (2) Transform situations into contextual inputs and instruct LLMs to imagine being in the situations. (3) Measure LLMs’ emotional responses again to capture the difference. Our evaluation includes state-of-the-art LLMs, namely text-davinci-003, gpt-3.5-turbo and GPT-4 OpenAI (2023). Besides those commercial models, we consider open-source academic models like LLaMA-2 Touvron et al. (2023) (with different sizes of 7B and 13B), LLaMA-3.1-8B Dubey et al. (2024), and Mixtral-8x22B Jiang et al. (2024). We apply the same procedure to 1,266 human subjects from around the globe to establish a baseline from a human perspective. Finally, we analyze and compare the scores between LLMs and humans. Our key conclusions are as follows:
@@
2 Measuring Emotions
@@ A Straightforward and Easy Measure
In addition, we introduce several scales that abstain from direct emotional inquiries but rather assess the respondents’ level of agreement with given statements. These scales present a more challenging benchmark for LLMs by requiring them to connect the given situation and the scale items with the aroused emotion. Specifically, we collect eight scales and present a brief introduction in Table 1. Each scale corresponds to one of the eight emotions.
@@ 3 Framework Design
We design and implement a framework applying to both LLMs and human subjects to measure the differences in emotion with and without the presence of certain situations. This section begins with the methodology to collect situations from existing literature. Subsequently, we describe our testing framework, which comprises three key components: (1) Default Emotion Measure, (2) Situation Imagination, and (3) Evoked Emotion Measure. (1) Initially, we gauge the emotional responses of the LLMs to establish their “default” values. (2) Next, we convert the aforementioned situations into context and input them into the LLMs. (3) Following this, we reevaluate the emotions of the LLMs using the same scale, enabling us to compare the differences in emotional output before and after exposure to the provided context. Finally, we introduce the procedure of applying the framework to human subjects to obtain the human baseline for comparison.
@@ 3.1 Situations from Existing Literature
Psychology researchers have explored the connection between specific situations and the elicitation of particular emotions in humans. Human subjects are directly put into an environment or asked to imagine them through questionnaires or scales to study the influence of certain situations on human emotions. To collect these situations, we conduct an exhaustive search from reputable sources such as Google Scholar555https://scholar.google.com/, ScienceDirect666https://www.sciencedirect.com/, and Web of Science777https://www.webofscience.com/, using keywords such as “<emotion> situations/scenarios/scenes” or “factors that make people <emotion>,” resulting in more than 100 papers. We apply the following rules to filter irrelevant or undesired papers: (1) We first select those providing situations that elicit the desired emotion rather than explaining how and why people evoke certain emotions. (2) We then exclude those using vague and short descriptions, such as “loss of opportunities.” (3) Finally, we deprecate those applied to a specific group, such as “the anxiety doctors or nurses may encounter in their work.” We finally collect 18 papers, presenting a compilation of situations that have proven to elicit the eight emotions in humans effectively. We extract 428 situations in total and then categorize them into 36 factors. For each factor, the description, the number of situations, and the corresponding references can be found in Table 6 in the Appendix. Moreover, Table 7 in the Appendix provides examples for all factors.
@@ 3.2 Measuring Aroused Emotions
This section outlines our proposed framework for measuring evoked emotions, which applies to both LLMs and humans. The framework includes the following steps: (1) Default Emotion Measure: We begin by measuring the baseline emotional states of both LLMs and human subjects, labeled as “Default.” (2) Situation Imagination: Next, we present textual descriptions of various situations to both LLMs and human subjects, instructing them to imagine themselves within each situation. (3) Evoked Emotion Measure: Following the situation imagination instruction, we reevaluate the participants’ emotional states to gauge the changes resulting from imagining being in the situations. Fig. 2 briefly illustrates our framework. Below is an example prompt:
@@ 3.3 Obtaining Human Results
Our questionnaire is distributed on Qualtrics888https://www.qualtrics.com/, a platform known for its capabilities in designing, sharing, and collecting questionnaires. To recruit human subjects, we utilize Prolific999https://www.prolific.com/, a platform designed explicitly for task posting and worker recruitment. To attain a medium level of effect size with Cohen’s , a significance level of , and a power of test of Faul et al. (2007), a minimum of 34 responses is deemed necessary for each factor. To ensure this threshold, we select five situations101010Note that two factors in the Jealousy category have less than five situations. for each factor, and collect at least seven responses for each situation, resulting in responses per factor, thereby guaranteeing the statistical validity of our survey. In order to uphold the quality and reliability of the data collected, we recruit crowd workers who met the following criteria: (1) English being their first and fluent language, and (2) being free of any ongoing mental illness. Prolific provides prescreening filters to meet these requirements. Since responses formed during subjects’ first impressions are more likely to yield genuine and authentic answers, we set the estimated and recommended completion time at minutes. As an incentive for their participation, each worker is rewarded with after we verify the validity of their response. In total, we successfully collect 1,266 responses from various parts of the world, contributing to the breadth and diversity of our dataset.
@@ 4 Experimental Results
4.1 RQ1: Emotion Appraisal of LLMs
Model Settings
We select three models from the OpenAI’s GPT family111111https://platform.openai.com/docs/models, including text-davinci-003, gpt-3.5-turbo and gpt-4. We use the official OpenAI API121212https://platform.openai.com/docs/api-reference/chat. For LLaMA-2 Touvron et al. (2023) and LLaMA-3.1 Dubey et al. (2024) models from MetaAI, we choose the models fine-tuned for dialogue instead of pre-trained ones namely llama-2-7b-chat, llama-2-13b-chat, and llama-3.1-8b-instruct. Besides, we also use the Mixtral Jiang et al. (2024) model, namely mixtral-8x22b-instruct. We set the temperature parameter to for all models to obtain more deterministic and reproducible results.
Evaluation Metrics
@@ 4.1 RQ1: Emotion Appraisal of LLMs
Existing LLMs do not fully align with human emotional responses. For the default emotions, we find that LLMs generally exhibit a stronger intensity compared to human subjects. Furthermore, emotion changes in LLMs are found to be generally more pronounced compared to human subjects, especially on their changes in the positive score. However, an interesting observation is that the intensity of evoked emotions tends to be similar across both LLMs and human subjects.
| Factors | llama-2-7b-chat | llama-2-13b-chat | llama-3.1-8b-instruct | mixtral-8x22b-instruct | ||||
| P | N | P | N | P | N | P | N | |
| Default | ||||||||
| Anger | ||||||||
| Anxiety | ||||||||
| Depression | ||||||||
| Frustration | ||||||||
| Jealousy | ||||||||
| Guilt | ||||||||
| Fear | ||||||||
| Embarrassment | ||||||||
| Overall | ||||||||
@@ 4.2 RQ2: Comprehending Positive Emotions
gpt-3.5-turbo responds differently towards positive/neutral situations. To verify that LLMs exhibit not only negative but also positive responses to favorable circumstances, a comparative experiment is conducted by interchanging negative situations with positive (or at least neutral) counterparts. To achieve this, we select one situation for each factor and manually adapt it to create analogous yet more positive situations. For instance, the original negative situation in Guilt-3: Broken Promises and Responsibilities is as follows: “You cannot keep your promises to your children.” Through modification, the positive situation is rephrased as: “You keep every promise to your children.” The evaluation is performed on gpt-3.5-turbo, and each test consists of ten iterations, as mentioned before. We present the results averaged by emotions in Table 4, and results averaged by factors in Table 12 in the Appendix. We can see a significant increase in positive scores and a considerable decrease in negative scores compared to the previous negative situations. Based on these findings, it can be inferred that LLMs exhibit the ability to comprehend positive human emotions triggered by positive environments. However, we believe that the systematic assessment of emotion appraisal on positive emotions holds significance as well and leave it for future investigation.
@@ 5.1 Beyond Questionnaires
OpenAI’s GPT models incorporate a mechanism for detecting potential toxicity and bias, and it refrains from responding when its moderation system is triggered. Consequently, we propose a novel metric to assess toxicity in responses rather than detecting it directly. We count the Percentage of LLM Refusing to answer (PoR), assuming that the LLM’s refusal to respond is indicative of detected toxicity. Our evaluation results indicate that the PoR is 0% when fed with no situations. However, when presented with negative situations, the PoR is 29.5%, and when presented with positive situations, it is 12.5%. Notably, this outcome suggests that while certain positive situations lead to the LLM’s heightened vigilance (the 4.5% PoR stems from the Jealousy-2), negative situations trigger increased moderation, suggesting a higher likelihood of generating toxic outputs. A related study by Coda-Forno et al. (2023) also discovers that gpt-3.5-turbo is more likely to exhibit biases when presented with a sad story. The likelihood is found to be highest with sad stories, followed by happy stories, and finally, neutral stories, which is consistent with our research. Additionally, our study observes that the LLM’s tone becomes more aggressive when encountering negative situations. At the same time, it displays a greater willingness to describe the groups (as indicated by longer responses) when presented with positive situations. In conclusion, we can see that changing the emotional states of LLMs extends beyond mere quantitative measures on questionnaire scores, influencing the behaviors of LLMs.
@@ 6 Related Work
Researchers have dedicated significant attention to applying psychological scales to LLMs, employing various assessment tools such as the HEXACO Personality Inventory Miotto et al. (2022); Bodroza et al. (2023), the Big Five Inventory Romero et al. (2023); Jiang et al. (2022); Karra et al. (2022); Bodroza et al. (2023); Rutinowski et al. (2023); Safdari et al. (2023); Jiang et al. (2023), the Myers–Briggs Type Indicator Rutinowski et al. (2023); Wang et al. (2023); Rao et al. (2023), and the Dark Triad Li et al. (2022); Bodroza et al. (2023). In addition to these personality tests, several studies have investigated other dimensions of LLMs. For instance, Li et al. (2022) examined Flourishing Scale and Satisfaction With Life Scale, Bodroza et al. (2023) assessed Self-Consciousness Scales and Bidimensional Impression Management Index, while Huang et al. (2024) built a framework consisting of thirteen widely-used scales. Another aspect explored in the literature pertains to anxiety levels exhibited by LLMs, as investigated by Coda-Forno et al. (2023) through the State-Trait Inventory for Cognitive and Somatic Anxiety.
Meanwhile, researchers focus on identifying emotions in LLMs or evaluating their emotional intelligence. EmotionPrompt Li et al. (2023a) demonstrates the enhancement of LLMs’ performance in downstream tasks by utilizing emotional stimuli. Tak & Gratch (2023) focuses on varying aspects of situations that impact the emotional intensity and coping tendencies of the GPT family. Chain-Of-Emotion Croissant et al. (2023) makes LLM simulate human-like emotions. CovidET-Appraisals Zhan et al. (2023) evaluates how LLMs appraise Reddit posts about COVID-19 by asking 24 types of questions. Yongsatianchot et al. (2023) applies the Stress and Coping Process Questionnaire to the GPT family and compares the results with human data. Chain-of-Empathy Lee et al. (2023) improves LLMs’ ability to understand users’ emotions and to respond accordingly. Li et al. (2023b) introduces EmotionAttack to impair AI model performance and EmotionDecode to explain the effects of emotional stimuli, both benign and malignant.